OpenStudy (anonymous):

A helicopter takes off from the ground 800 feet from the base of a building. It flies in a straight line directly to the top of the building. If the building is 160 feet tall, what angle did the helicopter take off from?

OpenStudy (anonymous):

\[\frac{180 \tan ^{-1}\left(\frac{160}{800}\right)}{\pi }=11.31 {}^{\circ} \]

OpenStudy (anonymous):

thank you I was really lost @robtobey

OpenStudy (anonymous):

Your welcome and thank you for the medals.

OpenStudy (anonymous):

no problem